OpenStudy (anonymous):

Select a tool or technique such as manipulatives or mental math to find A(7) when A(n) = A(n-1)+5 and A(1)=-20. A(7)=? HELP!

OpenStudy (anonymous):

@dan815 @pooja195

OpenStudy (anonymous):

Oh no! The question changed because I took too long! Ughhh. it is now the same except that A(1)= -25

OpenStudy (anonymous):

A(n)-A(n-1)=5 common difference d=5 A(1)=-25 an=a1+(n-1)d A(7)=-25+(7-1)*5=?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Yes it is 5! Thank you so much!!!

OpenStudy (anonymous):

correct.